OpenStudy (anonymous):

The ice skating rink charges an hourly fee for skating and $3 to rent skates for the day. Gillian rented skates and skated for 3 hours and was charged $21. Which equation represents the cost, c(x), of ice skating as a function of x, the number of hours of skating?

OpenStudy (anonymous):

A)c(x) = 3x + 3 B)c(x) = 6x + 3 C)c(x) = 7x + 3 D)c(x) = 8x + 3

OpenStudy (anonymous):

hi! try using any of them with the case given for Gillian, replacing the x by the hours he skated...for example using A: \[c(3)=3\times3+3=12\] so A is not the answer, we should get $21